40m
QIL
Cryo_Lab
CTN
SUS_Lab
TCS_Lab
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Tue May 3 21:05:23 2022, yuta, Update, BHD, AS path alignment, removing a lot of green stuff 7x
|
Wed May 4 13:09:42 2022, Anchal, Update, BHD, IMC table balanced again, IMC is locking, YARM is locking
|
Wed May 4 15:08:59 2022, Tega, Update, BHD, XARM alignment to get flashing
|
Wed May 4 18:52:43 2022, Tega, Update, BHD, POP beam too high at POP SM4 on the ITMX chamber
|
Thu May 5 14:46:22 2022, Anchal, Update, BHD, POP beam height lowered, POP_SM4 raised
|
Thu May 5 17:05:31 2022, Tega, Update, BHD, IMC & X/Y-arm alignment
|
Mon May 9 18:43:03 2022, Anchal, Update, BHD, ITMX table layout corrected
|
Mon May 9 18:49:05 2022, Tega, Update, BHD, Relocate green TRX and TRY PDs/cams/optics from PSL table to BS table 
|
Tue May 10 15:46:38 2022, JC, Update, BHD, Relocate green TRX and TRY components from PSL table to BS table 
|
Tue May 10 18:25:43 2022, Paco, Update, BHD, Green Transmission path
|
Tue May 10 16:31:14 2022, Tega, Update, BHD, ITMX optlev return beam steered to QPD
|
Fri May 6 13:48:34 2022, Anchal, Update, BHD, WFS loop instability fixed
|
Thu May 12 19:55:01 2022, Tega, Update, BHD, AS path alignment 
|
Thu May 12 20:11:18 2022, Anchal, Update, BHD, BHDBS Output beams steered out to ITMY table
|
Thu May 12 20:24:29 2022, yuta, Update, BHD, POX and POY investigation
|
Mon May 16 13:22:59 2022, yuta, Update, BHD, REFL and AS paths aligned at AP table 
|
Mon May 16 19:14:17 2022, Anchal, Update, BHD, Camera set on AS path and BHDBS output path
|
Tue May 17 18:43:38 2022, Anchal, Update, BHD, Placed SRM in ITMY Chamber
|
Wed May 18 08:30:29 2022, Paco, Update, BHD, SRM OpLev
|
Thu May 19 08:51:40 2022, Paco, Update, BHD, SRM OpLev
|
Wed May 18 09:02:52 2022, Anchal, Update, BHD, WFS1 PD centered
|
Wed May 18 17:23:15 2022, Anchal, Update, BHD, Placed PRM in BS Chamber
|
Thu May 19 18:56:08 2022, yuta, Update, BHD, Current OSEM sensor values with all the suspensions aligned
|
Thu May 19 19:05:59 2022, Anchal, Update, BHD, BS Chamber all work finished, BHD path setup
|
Fri May 20 20:03:48 2022, Paco, Update, BHD, ITMY chamber work finished - LO and AS overlapped 
|
Mon May 23 13:16:59 2022, Paco, Update, BHD, End of vent - checks  
|
Tue May 24 15:21:13 2022, Anchal, Update, BHD, Freeswing tests of new SOS started
|
|
Message ID: 16843
Entry time: Tue May 10 16:31:14 2022
In reply to: 16837
|
Author: |
Tega |
Type: |
Update |
Category: |
BHD |
Subject: |
ITMX optlev return beam steered to QPD |
|
|
Followed the steps below to complete the ITMX optlev installation. The ITMX optlev return beam now reaches its QPD without being blocked by the input steering mirror.
Although, I centered the ITMX optlev readout, this was not done when the XARM flashing is maximized bcos the IMC chamber was being worked on, so this should be done later when the IR beam is back.
Quote: |
Next steps:
- The Oplev beam paths need to be adjusted.
- The ongoing beam steering mirror is blocking the returning beam, so the ongoing path needs to be changed.
- First setup two irises to save ingoing path.
- Then make space for the returning beam by changing the steering mirror positions.
- Then recover the ingoing path to the center of irises.
- Steer the returning beam to the QPD.
- Then maximize the flashing on XARM and center the oplev to save this position.
|
|